OpenStudy (anonymous):

Check my answers again Lol please Find the number of real number solutions for the equation. x^2-15=0 1 0 2 - my answer Can not be determined from the graph

OpenStudy (dinamix):

are u sure its x^2-15=0 o.O ?

OpenStudy (anonymous):

Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. A rocket is launched from atop a 101-foot cliff with an initial velocity of 116 ft/s. a. Substitute the values into the vertical motion formula h = -16t^2 + vt + c. Let h = 0 b. Use the quadratic formula find out how long the rocket will take to hit the ground after it is launched. Round to the nearest tenth of a second. A) 0 = -16t^2+101t+116; 8s B) 0 = -16t^2+116t+101; 0.8s - my answer C) 0 = -16t^2+101t+116; 0.8s D) 0 = -16t^2+116t+101; 8s

OpenStudy (phi):

\[ x^2-15=0 \\ x^2= 15 \\ x= \pm \sqrt{15} \] you would not get real number solutions if you had to take the square root of -15 but here, you have two real numbers.

OpenStudy (anonymous):

wait a min let me recheck

OpenStudy (anonymous):

Yeah @dinamix

OpenStudy (dinamix):

cuz i find answer like @phi

OpenStudy (phi):

you have the correct equation, but if we used t=1 (which is close to t=0.8) we get -16+116+101 which is not close to zero (we want height = 0 when on the ground)

OpenStudy (anonymous):

I don't get it, are my answers wrong..

OpenStudy (anonymous):

can you help me @ASAAD123

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

and one more question Use any method to solve the equation. If necessary, round to the nearest hundredth. 7x^2-16x=8 0.42, –2.71- my answer 35.43, –33.14 5.42, –2.95 2.71, –0.42

OpenStudy (anonymous):

the first one is correct. The equation is correct for the second one. The solution isn't. The third one is wrong

OpenStudy (anonymous):

I wasn't sure.. for the second one I was like either B or C are both wrong.. if not B then C, not sure tho @peachpi

OpenStudy (anonymous):

the equation you selected is the right one, but 0.8 isn't a solution to it. Check to see that you plugged the numbers into the quadratic equation correctly.

OpenStudy (anonymous):

the third one is it D?

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

Can someone helpp meee

OpenStudy (anonymous):

which question?

OpenStudy (anonymous):

and one more question Use any method to solve the equation. If necessary, round to the nearest hundredth. 7x^2-16x=8 0.42, –2.71- my answer 35.43, –33.14 5.42, –2.95 2.71, –0.42 - my answer

OpenStudy (anonymous):

D

OpenStudy (anonymous):

not A

OpenStudy (anonymous):

so we are trying to find solution to \[7x^{2}-16x=8\] ?

OpenStudy (anonymous):

so which method did you use to solve this question?

OpenStudy (anonymous):

you could use, graphical methods, quadratic formula, etc..

OpenStudy (anonymous):

yes D